Below is a partially complete probability model. 
Enter the probability for the final outcome. 

Outcome     | 0   |  2.5 |    4 | 5.5 | 
Probability |0.25 | 0.25 | 0.15 |  ?  |  

Find the expected value of the probability model.

Answer by Edwin McCravy(20064)

Below is a partially complete probability model. 
Enter the probability for the final outcome. 

Outcome     | 0   |  2.5 |    4 | 5.5 | 
Probability |0.25 | 0.25 | 0.15 |  ?  |  

------------------------------------------

All the probabilities of a probability model must have sum 1

0.25 + 0.25 + 0.15 + ? = 1.0
              0.65 + ? = 1.0
                     ? = 1.00 - 0.65
                     ? = 0.35

Outcome     | 0   |  2.5 |    4 |  5.5 | 
Probability |0.25 | 0.25 | 0.15 | 0.35 |

-------------------------------------------------

Find the expected value of the probability model.

-------------------------------------------------

Multiply the value of each outcome by its probability and add them:

E = (0)(0.25)+(2.5)(0.25)+(4)(0.15)+(5.5)(0.35) = 3.15

Edwin